Three different plans for financing an $8,400,000 corporation are under consideration by its organizers. Under each of the following plans, the securities will be issued at their par or face amount, and the income tax rate is estimated at 40% of income:

Plan 1 Plan 2 Plan 3
10% bonds _ _ $4,200,000
Preferred 5% stock, $80 par _ $4,200,000 2,100,000
Common stock, $8.4 par $8,400,000 4,200,000 2,100,000
Total $ 8,400,000 $ 8,400,000 $ 8,400,000

Required:

1. Determine the earnings per share of common stock for each plan, assuming that the income before bond interest and income tax is $16,800,000. Enter answers in dollars and cents, rounding to the nearest cent.

Earnings Per Share on Common Stock
Plan 1 $
Plan 2 $
Plan 3 $

2. Determine the earnings per share of common stock for each plan, assuming that the income before bond interest and income tax is $7,980,000. Enter answers in dollars and cents, rounding to the nearest cent.

Earnings Per Share on Common Stock
Plan 1 $
Plan 2 $
Plan 3 $
Answer #1

Answer-1)-

EARNING PER SHARE ON COMMON STOCK
Particluars Plan 1 Plan 2 Plan 3
$ $ $
Income before interest & taxes 16800000 16800000 16800000
Less:- Interest on bonds 0 0 4200000*10%= 420000
Income before taxes 16800000 16800000 16380000
Income Taxes 40% 16800000*40%=6720000 16800000*40%= 6720000 16380000*40%= 6552000
Income after taxes(A) 10080000 10080000 9828000
Less- Preferred dividends 0 4200000*5%= 210000 2100000*5%= 105000
Earning for common stockholders 10080000 9870000 9723000
No of shares (B) 1000000 500000 250000
Earning per share C=A/B 13.81 20.16 38.89

2)-

EARNING PER SHARE ON COMMON STOCK
Particluars Plan 1 Plan 2 Plan 3
$ $ $
Income before interest & taxes 7980000 7980000 7980000
Less:- Interest on bonds 0 0 4200000*10%= 420000
Income before taxes 7980000 7980000 7560000
Income Taxes 40% 7980000*40%=3192000 7980000*40%= 3192000 7560000*40%= 3024000
Income after taxes(A) 4788000 4788000 4536000
Less- Preferred dividends 0 4200000*5%= 210000 2100000*5%= 105000
Earning for common stockholders 4788000 4578000 4431000
No of shares (B) 1000000 500000 250000
Earning per share C=A/B 6.56 9.58 17.72